# DialectGen: Benchmarking and Improving Dialect Robustness in Multimodal Generation

### v1

@@ Abstract
Contact languages like English exhibit rich regional variations in the form of dialects, which are often used by dialect speakers interacting with generative models. However, can multimodal generative models effectively produce content given dialectal textual input? In this work, we study this question by constructing a new large-scale benchmark spanning six common English dialects. We work with dialect speakers to collect and verify over 4,200 unique prompts and evaluate on 17 image and video generative models. Our automatic and human evaluation results show that current state-of-the-art multimodal generative models exhibit 32.26% to 48.17% performance degradation when a single dialect word is used in the prompt. Common mitigation methods such as fine-tuning and prompt rewriting can only improve dialect performance by small margins (< 7%), while potentially incurring significant performance degradation in Standard American English (SAE). To this end, we design a general encoder-based mitigation strategy for multimodal generative models. Our method teaches the model to recognize new dialect features while preserving SAE performance. Experiments on models such as Stable Diffusion 1.5 show that our method is able to simultaneously raise performance on five dialects to be on par with SAE (+34.4%), while incurring near-zero cost to SAE performance. Code and data at: dialectgen.github.io.
@@ 1 Introduction
Linguists have defined over 160 dialects (Aeni et al., 2021) within the English language, with three out of four English speakers having a dialect background other than Standard American or British English (Crystal, 2003). Despite this rich diversity, current pre-training paradigms employ content filters that can exclude data involving lower-resource English dialects other than Standard American and British English (Gururangan et al., 2022), reducing the effectiveness of pretrained models on inputs from other dialects (Lee et al., 2023). Prior works have shown significant allocational harms toward dialect speakers caused by such dialect performance discrepancies in machine learning applications (Hovy & Spruit, 2016; Bender et al., 2021), making the observation of similar performance trends in multimodal generative models an alarming sign.
As shown in Figure˜1, while current multimodal generative models can accurately generate high quality image and video content given Standard American English (SAE) prompts (left); they fail in various manners when provided with semantically equivalent prompts containing a single synonymous dialect word (right). Stable Diffusion 3.5 Large (Esser et al., 2024) fails to generate "ang pow", which is commonly used in Singaporean English to mean "red packet", and FLUX.1 [dev] (Black Forest Labs, 2024) fails to generate "brinjal", which is synonymous with "eggplant" in Indian English. Furthermore, when the dialect lexeme is polysemous, i.e., has an alternative meaning in SAE, models tend to always generate content that align with the SAE meaning, even when the context makes such interpretation highly improbable. For example, DALL-E Mini (Dayma et al., 2021) generations of "A man driving his whip" fail to capture the correct meaning of "whip" as "car" in African American English, given clear context indications. Similar failure modes are observed in text-to-video generative models: Wan 2.1 (Wang et al., 2025) fails to correctly render "carnal", which refers to "brother" in Chicano English.
@@ 2 Related Works
Linguists define dialects as regional variations of a language distinguished by unique features in lexicon, phonology, and grammar from each other, together constituting a single language (Hudson, 1996; Chambers & Trudgill, 1998; Fromkin et al., 1998; Nerbonne, 2009; Wardhaugh & Fuller, 2021). English, like any other language, is subject to such variations. However, most dataset resources and pre-training paradigms focus only on Standard American and British English (Gururangan et al., 2022), leading to dialect robustness issues and performance gaps in downstream machine learning applications. Previous works have analyzed and explored such dialectal performance gaps in NLP tasks like QA (Ziems et al., 2023), NLI (Ziems et al., 2022), dependency parsing, and POS tagging (Blodgett et al., 2018; Jørgensen et al., 2015). Recent works have also noticed the impact of dialect variations on text-to-image generation (Lee et al., 2023; Wan et al., 2024). Along this line of research, we create the first large-scale benchmark of dialect robustness in multimodal generation, evaluating both text-to-image and text-to-video generative models on inputs across six different dialects.
@@ 3.1 Dataset Construction
To select dialect features for our benchmark dataset, we first gather dialect lexemes along with their dictionary definitions and example usages from publicly available regional English dictionaries including The Oxford Regional English Dictionary (Gates et al., 2023), Dictionary of American regional english (Cassidy et al., 1985), A dictionary of Singlish and Singapore English (Lee, 2004), Dictionary of Indian English (Subhash, 2020), and The Oxford Dictionary of African American English (Heinmiller, 2023). We collect a total of 1126 dialect lexemes for initial processing.
@@ 4 Experiments
Concise Prompts
T2I Models
Stable Diffusion 1.4 28.19 26.7 10.35 20.67 9.64 34.94 41.27 26.96
@@
Detailed Prompts
T2I Models
Stable Diffusion 1.4 14.33 15.93 5.16 11.65 4.37 17.35 29.23 17.03
@@ 4.2 Benchmark Experiments
Looking at individual model performances, we observe that among text-to-video generative models: Wan 2.1 (Wang et al., 2025) and CogVideoX (Yang et al., 2024) exhibit the largest overall performance drops while Cosmos-1 (Agarwal et al., 2025) is the most robust. While for text-to-image generative models, DALL-E 2 (Ramesh et al., 2022) and Flux.1 [dev] (Black Forest Labs, 2024) exhibit the largest overall performance drops while DALL-E 3 (Betker et al., 2023) (w/ Prompt Rewrite) and gpt-image-1 (4o Image Generation) (OpenAI, 2025) are the most robust.
@@ 5.1 Baseline Methods
  • Prompt Translate We use general-purpose LLMs (Grattafiori et al., 2024; OpenAI, 2025) to translate all prompts to SAE before passing them to the generative model.
@@ 5.3.2 Ablation Study
To solve this issue, simply maximizing cosine similarity between the target text encoder’s text embeddings and the corresponding text/image embeddings from the frozen text/image encoder (denoted as Text/Image Cosine Reg.), which is computed over the same caption-image pairs used in our KL regularization, does not effectively recover the base model’s SAE MSCOCO and polysemy performance. In contrast, adding our KL regularization loss ($\mathcal{L}_{\text{KL}}$) improves both metrics while preserving dialect gains.
@@ 6 Limitations
Our study focuses on the lexical variations that characterize dialects, motivated by the empirical observation that such variations exert much greater influence on multimodal generative model performance than grammatical variations (see Appendix˜G). Furthermore, grammatical variation has already been the subject of extensive investigation in text-only contexts (Hudson, 1996; Chambers & Trudgill, 1998; Fromkin et al., 1998; Nerbonne, 2009; Wardhaugh & Fuller, 2021). These considerations jointly motivate our decision to prioritize the evaluation of lexical dialect variation, which appears especially consequential in the multimodal generative setting. Furthermore, our evaluation of text-image alignment utilizes reference-free metrics, namely VQAScore (Lin et al., 2024) and CLIPScore (Hessel et al., 2021). We recognize that these pretrained vision-language models are not perfect. To address this potential weakness, we conducted a thorough human evaluation and found very high statistical correlation between our automatic metrics and human judgment (Pearson correlation coefficient $r$ = 0.968 for VQAScore and $r$ = 0.924 for CLIPScore). Therefore, while acknowledging the imperfections of automated metrics, this high degree of human correlation provides strong evidence for the validity of our evaluation metrics and associated analysis conclusions.
@@ Appendix F Mitigation Results on Stable Diffusion XL
Concise Prompts
T2I Models
Stable Diffusion 1.4 60.66 76.47 71.46 79.08 51.31 78.86 47.5 80.88 57.64 78.92
@@
Detailed Prompts
T2I Models
Stable Diffusion 1.4 70.07 79.31 74.19 77.58 65.24 78.94 56.99 80.53 63.87 76.98
@@
Concise Prompts
T2I Models
Stable Diffusion 1.4 25.46 27.83 28.65 29.79 24.68 28.34 24.34 29.38 25.97 28.64
@@
Detailed Prompts
T2I Models
Stable Diffusion 1.4 27.98 28.84 29.59 30.23 28.61 30.1 26.79 29.83 28.12 29.78
@@
Appendix H Performance by Dialect
@@ Appendix I Use of AI tools
Appendix J Future Work
@@
Many existing studies rely on the accurate semantic understanding and high-fidelity generation capabilities of multimodal text-to-image and text-to-video generative models Zhang et al. (2023); Wallace et al. (2024); Zhou et al. (2025). It would be interesting to investigate the downstream research impacts of dialectal performance gaps on these works as well as downstream societal impacts to dialect speaker user groups.

### v2

@@ Abstract
Contact languages like English exhibit rich regional variations in the form of dialects, which are often used by dialect speakers interacting with generative models. However, can multimodal generative models effectively produce content given dialectal textual input? In this work, we study this question by constructing a new large-scale benchmark spanning six common English dialects. We work with dialect speakers to collect and verify over 4,200 unique prompts and evaluate on 17 image and video generative models. Our automatic and human evaluation results show that current state-of-the-art multimodal generative models exhibit 32.26% to 48.17% performance degradation when a single dialect word is used in the prompt. Common mitigation methods such as fine-tuning and prompt rewriting can only improve dialect performance by small margins (< 7%), while potentially incurring significant performance degradation in Standard American English (SAE). To this end, we design a general encoder-based mitigation strategy for multimodal generative models. Our method teaches the model to recognize new dialect features while preserving SAE performance. Experiments on models such as Stable Diffusion 1.5 show that our method is able to simultaneously raise performance on five dialects to be on par with SAE (+34.4%), while incurring near-zero cost to SAE performance.
@@ 1 Introduction
Linguists have defined over 160 dialects (Aeni et al., 2021) within the English language, with three out of four English speakers having a dialect background other than Standard American or British English (Crystal, 2003). Despite this rich diversity, current pre-training paradigms employ content filters that can exclude data involving lower-resource English dialects other than Standard American and British English (Gururangan et al., 2022), reducing the effectiveness of pretrained models on inputs from other dialects (Lee et al., 2023). Prior works have shown significant allocational harms toward dialect speakers caused by such dialect performance discrepancies in machine learning applications (Hovy and Spruit, 2016; Bender et al., 2021), making the observation of similar performance trends in multimodal generative models an alarming sign.
As shown in Figure˜1, while current multimodal generative models can accurately generate high quality image and video content given Standard American English (SAE) prompts (left); they fail in various manners when provided with semantically equivalent prompts containing a single synonymous dialect word (right). Stable Diffusion 3.5 Large (Esser et al., 2024) fails to generate "ang pow", which is commonly used in Singaporean English to mean "red packet", and FLUX.1 [dev] (Black Forest Labs, 2024) fails to generate "brinjal", which is synonymous with "eggplant" in Indian English. Furthermore, when the dialect lexeme is polysemous, i.e., has an alternative meaning in SAE, models tend to always generate content that align with the SAE meaning, even when the context makes such interpretation highly improbable. For example, DALL-E Mini (Dayma_DALL·E_Mini_2021) generations of "A man driving his whip" fail to capture the correct meaning of "whip" as "car" in African American English, given clear context indications. Similar failure modes are observed in text-to-video generative models: Wan 2.1 (Wang et al., 2025) fails to correctly render "carnal", which refers to "brother" in Chicano English.
@@ 2 Related Works
Linguists define dialects as regional variations of a language distinguished by unique features in lexicon, phonology, and grammar from each other, together constituting a single language (Hudson, 1996; Chambers and Trudgill, 1998; Fromkin et al., 1998; Nerbonne, 2009; Wardhaugh and Fuller, 2021). English, like any other language, is subject to such variations. However, most dataset resources and pre-training paradigms focus only on Standard American and British English (Gururangan et al., 2022), leading to dialect robustness issues and performance gaps in downstream machine learning applications. Previous works have analyzed and explored such dialectal performance gaps in NLP tasks like QA (Ziems et al., 2023), NLI (Ziems et al., 2022), dependency parsing, and POS tagging (Blodgett et al., 2018; Jørgensen et al., 2015). Recent works have also noticed the impact of dialect variations on text-to-image generation (Lee et al., 2023; Wan et al., 2024). Along this line of research, we create the first large-scale benchmark of dialect robustness in multimodal generation, evaluating both text-to-image and text-to-video generative models on inputs across six different dialects.
@@ 3.1 Dataset Construction
To select dialect features for our benchmark dataset, we first gather dialect lexemes along with their dictionary definitions and example usages from publicly available dictionaries including The Oxford English Dictionary Regional English Database  [^1] https://www.oed.com/search/advanced/Entries?regional, Dictionary of American Regional English (Cassidy, 1985), A Dictionary of Singlish and Singapore English (Lee, 2004), Dictionary of Indian English (Subhash, 2020), and The Oxford Dictionary of African American English (Heinmiller, 2023). We collect a total of 1126 dialect lexemes for initial processing.
@@ 4 Experiments
Concise Prompts T2I Models Stable Diffusion 1.4 28.19 26.7 10.35 20.67 9.64 34.94 41.27 26.96
@@
Detailed Prompts T2I Models Stable Diffusion 1.4 14.33 15.93 5.16 11.65 4.37 17.35 29.23 17.03
@@ 4.2 Benchmark Experiments
Looking at individual model performances, we observe that among text-to-video generative models: Wan 2.1 (Wang et al., 2025) and CogVideoX (Yang et al., 2024) exhibit the largest overall performance drops while Cosmos-1 (Agarwal et al., 2025) is the most robust. While for text-to-image generative models, DALL-E 2 (Ramesh et al., 2022) and Flux.1 [dev] (Black Forest Labs, 2024) exhibit the largest overall performance drops while DALL-E 3 (Betker et al., 2023) (w/ Prompt Rewrite) and gpt-image-1 (4o Image Generation) (OpenAI, 2025b) are the most robust.
@@ 5.1 Baseline Methods
  • Prompt Translate We use general-purpose LLMs (Grattafiori et al., 2024; OpenAI, 2025a) to translate all prompts to SAE before passing them to the generative model.
@@ 5.3.2 Ablation Study
Simply maximizing cosine similarity between the target text encoder’s text embeddings and the corresponding text/image embeddings from the frozen text/image encoder (denoted as Text/Image Cosine Reg.), which is computed over the same caption-image pairs used in our KL regularization, does not effectively recover the base model’s SAE MSCOCO and polysemy performance. In contrast, adding our KL regularization loss ($\mathcal{L}_{\text{KL}}$) improves both metrics while preserving dialect gains.
@@ 6 Limitations
Our study focuses on the lexical variations that characterize dialects, motivated by the empirical observation that such variations exert much greater influence on multimodal generative model performance than grammatical variations (see Appendix˜G). Furthermore, grammatical variation has already been the subject of extensive investigation in text-only contexts (Hudson, 1996; Chambers and Trudgill, 1998; Fromkin et al., 1998; Nerbonne, 2009; Wardhaugh and Fuller, 2021). These considerations jointly motivate our decision to prioritize the evaluation of lexical dialect variation, which appears especially consequential in the multimodal generative setting. Furthermore, our evaluation of text-image alignment utilizes reference-free metrics, namely VQAScore (Lin et al., 2024) and CLIPScore (Hessel et al., 2021). We recognize that these pretrained vision-language models are not perfect. To address this potential weakness, we conducted a thorough human evaluation and found very high statistical correlation between our automatic metrics and human judgment (Pearson correlation coefficient $r$ = 0.968 for VQAScore and $r$ = 0.924 for CLIPScore). Therefore, while acknowledging the imperfections of automated metrics, this high degree of human correlation provides strong evidence for the validity of our evaluation metrics and associated analysis conclusions.
@@ Appendix F Mitigation Results on Stable Diffusion XL
Appendix H Compositional Impact of Multiple Dialect Lexemes on Multimodal Generative Models
In this section we share results of a quantitative experiment on 100 newly annotated SAE-dialect prompt pairs from the Indian English setting Table˜9. Each prompt contains 2-3 SAE lexemes with dialect word equivalents. In the Single Dialect Word setting, we replace one of them with its dialect word equivalent, and in the Multi Dialect Word setting we replace 2-3 words. Finally, we evaluate each model’s generation quality with the VQAScore metric.
In line with our expectations, models experience more significant performance degradation when more dialectal lexemes are introduced in the prompt. These results highlight the challenging nature of the dialect robustness problem in multimodal generation. It further hints at the compositional impact of the issue, where dialect understanding errors become more severe when more dialectal lexicons are involved. We look forward to further exploring this topic in our future works.
Table 9: Quantitative Effects of Single vs. Multiple Dialectal Lexeme Variations on Multimodal Generation, measured in VQAScore. We evaluate three text-to-image generative models on 100 newly annotated Indian English SAE–dialect prompt pairs. The Single Dialect Lexeme setting replaces one SAE word with its dialectal equivalent, while the Multi Dialect Lexeme setting replaces 2–3 words. Values in parentheses indicate the percentage performance drop relative to SAE prompts.
Model SAE Prompt Performance (%) Dialectal Prompt Performance (%)
Single Dialect Lexeme Multi Dialect Lexeme
DALL·E Mini 72.39 48.65 (-32.79) 32.47 (-55.15)
FLUX.1 dev 83.04 54.81 (-34.00) 38.90 (-46.26)
Stable Diffusion 3.5 Large 88.73 56.21 (-36.65) 37.62 (-57.60)
@@ Appendix H Compositional Impact of Multiple Dialect Lexemes on Multimodal Generative Models
Concise Prompts T2I Models Stable Diffusion 1.4 60.66 76.47 71.46 79.08 51.31 78.86 47.5 80.88 57.64 78.92
@@
Detailed Prompts T2I Models Stable Diffusion 1.4 70.07 79.31 74.19 77.58 65.24 78.94 56.99 80.53 63.87 76.98
@@
Concise Prompts T2I Models Stable Diffusion 1.4 25.46 27.83 28.65 29.79 24.68 28.34 24.34 29.38 25.97 28.64
@@
Detailed Prompts T2I Models Stable Diffusion 1.4 27.98 28.84 29.59 30.23 28.61 30.1 26.79 29.83 28.12 29.78
@@
Appendix I Performance by Dialect
@@
Appendix K Future Work
@@
Many existing research studies Zhang et al. (2023); Wallace et al. (2024); Geng et al. (2024); Zhou et al. (2025); Li et al. (2025) and industrial systems rely on the accurate semantic understanding and high-fidelity generation capabilities of multimodal text-to-image and text-to-video generative models. It would be interesting to investigate the downstream research impacts of dialectal performance gaps on these works as well as downstream societal impacts to dialect speaker user groups.